OpenStudy (vortish):

complet the following parts y -x=-2 (a) write the equation in slope intercept form (b) give the slope and y intercept of the line

OpenStudy (anonymous):

slope intercept form is y=mx+b m = is the slope b = y intercept so solve for y y -x=-2 y = x - 2

OpenStudy (anonymous):

y-x=-2 when ever you want to find the y value x will be zero witch is y-0=-2 and the opposite for x like 0-x=-2 see

OpenStudy (anonymous):

now you want to divide y and x for two what do you do?

OpenStudy (vortish):

i am having a hard time decifering your methology

OpenStudy (anonymous):

ao waite

OpenStudy (vortish):

if i am following 0-x=-2 x= 2 and the same for y-x=-2 x = 2 as well or i did some thing worng

OpenStudy (anonymous):

ao you did something wrong let me explain it in numbers instead of words

OpenStudy (anonymous):

you want Y to be al alone so what do you do?

OpenStudy (anonymous):

y-x=2 you will add x for both side if the equation like this y-x=2 y-x=2 +2 +2 y=x-2

OpenStudy (anonymous):

technically u just make the x positive and take it to the other side of the sign(=).